# ProDAG: Projection-Induced Variational Inference for Directed Acyclic Graphs

### v4

@@ Appendix B Gradients
The acyclicity constraints forces $\|\omega_{jk}^{h}\|_{2}=0$ for some $j$ and $k$, meaning the whole vector $\omega_{jk}^{h}=0$. Likewise, the sparsity constraint, which represents a group lasso regularizer (Yuan & Lin, 2006), is also known to set some $\omega_{jk}^{h}$ to zero. Those surviving vectors that are not set to zero are shrunk towards zero by an amount $\|\omega_{jk}^{h}\|_{2}/\|\tilde{\omega}_{jk}^{h}\|_{2}$. The combined effect of these two constraints gives that
@@
Appendix D Implementation
ProDAG uses a multivariate Gaussian prior on $\tilde{W}$ with independent elements, each having mean zero and variance one. The variational posterior on $\tilde{W}$ is also taken as a multivariate Gaussian with independent elements, where the means and variances are initialized at their prior values. We use the Adam optimizer (Kingma & Ba, 2015) with a learning rate of 0.1 to minimize the variational objective function. The posterior variances are held positive during optimization using a softplus transform. At each Adam iteration, 100 samples of $\tilde{W}$ are drawn from the posterior to estimate the objective function. To project these sampled $\tilde{W}$ and obtain $W$, we solve the projection using gradient descent with learning rates of $1/p$ and $0.25/p$ for the linear and nonlinear settings, respectively. Similar to DAGMA, the resulting $W$ can contain small but nonzero elements that violate acyclicity. These small nonzeros occur because gradient descent and related first-order methods seldom produce machine-precision zeros. To that end, we follow standard practice and set small elements of $W$ to zero with a threshold of 0.1. At inference time, this threshold can be set such that $W$ is guaranteed acyclic (see, e.g., Ng et al., 2020).
For the benchmark methods, we use open-source Python or Julia implementations:
We use 50 particles for DiBS’s particle variational inference in the synthetic data examples and 10 particles in the real data example. More particles exceeded available memory.
@@ Appendix D Implementation
The grids are chosen so that the sparsity level matches the true sparsity level for some value between $\lambda^{\mathrm{min}}$ and $\lambda^{\mathrm{max}}$. For the real dataset, we choose $\lambda$ from the grid according to the sparsity of the consensus graph. For the synthetic datasets, we choose $\lambda$ using a separate validation set of size $\lfloor 0.1n\rceil$. We cannot extract the weighted adjacency matrices or neural networks from BayesDAG (only binary adjacency matrices), so we choose its $\lambda$ by revealing the true sparsity level.
In the nonlinear setting, the neural networks for ProDAG and DiBS consist of a single hidden layer with 10 neurons that use ReLU activation functions. DAGMA employs the same architecture but the neurons use sigmoid activations (its implementation does not support custom activation functions). The network for BayesDAG consists of two hidden layers, each with 128 neurons (its implementation does not support custom numbers of layers or neurons) that use ReLU activations.
The experiments are run on a Linux workstation with an AMD Ryzen Threadripper PRO 5995WX CPU, 256GB RAM, and 2 × NVIDIA GeForce RTX 4090 GPU. Individual experiments are run on either a single GPU or a single CPU core.
Appendix E Simulation Design
The synthetic datasets generated from linear DAGs are constructed as follows. We first randomly sample an Erdős–Rényi graph with $p$ nodes and $s$ edges. The edges of this undirected graph are then oriented according to a randomly generated topological ordering. Each directed edge in the graph is then weighted by sampling weights uniformly on $[-0.7,-0.3]\cup[0.3,0.7]$, yielding a weighted adjacency matrix $W$. We then draw iid noise $\varepsilon_{1},\dots,\varepsilon_{n}$ and finally generate the variables as
@@ Appendix E Simulation Design
For synthetic datasets generated from nonlinear DAGs, we first sample an Erdős–Rényi graph and orientate it according to a random topological ordering. We then generate a feedforward neural network $f_{k}$ with a single hidden layer of 10 neurons for $k=1,\dots,p$. The weights of $f_{k}$ are sampled uniformly from $[-0.7,-0.3]\cup[0.3,0.7]$, with hidden layer weights of node $k$’s non-parents set to zero. The activation is a ReLU. The noise $\varepsilon_{1},\dots,\varepsilon_{n}$ is drawn iid and the variables are generated as
@@
Appendix F Additional Experiments
@@
Figure 7: Performance on synthetic datasets generated from linear scale-free DAGs with $p=20$ nodes, $s=40$ edges, and Gaussian noise. The averages (solid points) and standard errors (error bars) are measured over 10 independently and identically generated datasets.
We investigate the robustness of ProDAG to several alternative simulation designs. Figure 5 reports results on datasets simulated from dense graphs containing twice as many edges as used in the main experiments. Figure 6 reports the results from a non-Gaussian noise distribution, specifically a $\operatorname{Gumbel}(0,1)$ distribution. Finally, Figure 7 reports results on scale-free graphs instead of Erdős–Rényi graphs. Under all three simulation designs, the relative performance of ProDAG is roughly similar to that in the main experiments, indicating that ProDAG is reliable across a broad range of scenarios.

### v5

@@ Appendix D Proof of Proposition 1
The acyclicity constraint forces $\|\omega_{jk}^{h}\|_{2}=0$ for some $j$ and $k$, meaning the whole vector $\omega_{jk}^{h}=0$. Likewise, the sparsity constraint, which represents a group lasso regularizer (Yuan and Lin, 2006), is also known to set some $\omega_{jk}^{h}$ to zero. Those surviving vectors that are not set to zero are shrunk towards zero by a factor $\|\omega_{jk}^{h}\|_{2}/\|\tilde{\omega}_{jk}^{h}\|_{2}$. The combined effect of these two constraints gives
@@
Appendix E Implementation
ProDAG uses a multivariate Gaussian prior on $\tilde{W}$ with independent elements, each having mean zero and variance one. The variational posterior on $\tilde{W}$ is also taken as a multivariate Gaussian with independent elements, where the means and variances are initialized at their prior values. We use the Adam optimizer (Kingma and Ba, 2015) with a learning rate of 0.1 to minimize the variational objective function. The posterior variances are held positive during optimization using a softplus transform. At each Adam iteration, 100 samples of $\tilde{W}$ are drawn from the posterior to estimate the objective function. To project these sampled $\tilde{W}$ and obtain $W$, we solve the projection using gradient descent with learning rates of $1/p$ and $0.25/p$ for the linear and nonlinear settings, respectively. Similar to DAGMA, the resulting $W$ can contain small but nonzero elements that violate acyclicity. These small nonzeros occur because gradient descent seldom produces machine-precision zeros. To that end, we follow standard practice and set small elements of $W$ to zero with a threshold of 0.1. At inference time, this threshold can be set such that $W$ is guaranteed acyclic (see, e.g., Ng et al., 2020).
Our implementation of ProDAG uses the machine learning library Flux (Innes et al., 2018) and performs projections using parallel GPU (batched CUDA) implementations of the projection algorithms.
For the benchmark methods, we use the respective authors’ open-source Python implementations:
We use 50 particles for DiBS’s particle variational inference. More particles exceed available memory.
@@ Appendix E Implementation
The above grids are chosen so that the sparsity level matches the true sparsity level for some value between $\lambda^{\mathrm{min}}$ and $\lambda^{\mathrm{max}}$. We choose the specific value of $\lambda$ using a separate validation set of size $\lfloor 0.1n\rceil$. We cannot extract the weighted adjacency matrices or neural networks from BayesDAG (only the binary adjacency matrices), so we choose its $\lambda$ so that it is closest to the true sparsity level.
In the nonlinear setting, the neural networks for ProDAG and DiBS consist of a single hidden layer with 10 neurons that use ReLU activation functions. DAGMA employs the same architecture but the neurons use sigmoid activations as its implementation does not support custom activation functions. The network for BayesDAG consists of two hidden layers, each with 128 neurons that use ReLU activations. BayesDAG’s implementation does not support custom numbers of layers or neurons.
The experiments are run on a Linux workstation with an AMD Ryzen Threadripper PRO 5995WX CPU, 256GB RAM, and 2 × NVIDIA GeForce RTX 4090 GPUs. Each method is allocated either a single GPU or a single CPU core, with the experiments run in parallel across GPUs/cores.
Appendix F Simulation design
The synthetic datasets generated from linear DAGs are constructed as follows. We first randomly sample an Erdős–Rényi or scale-free graph with $p$ nodes and $s$ edges. The edges of this undirected graph are then oriented according to a randomly generated topological ordering. Each directed edge in the graph is then weighted by sampling weights uniformly on $[-0.7,-0.3]\cup[0.3,0.7]$, yielding a weighted adjacency matrix $W$. We then draw iid noise $\varepsilon_{1},\dots,\varepsilon_{n}$ and generate the variables as
@@ Appendix F Simulation design
For synthetic datasets generated from nonlinear DAGs, we likewise sample an Erdős–Rényi or scale-free graph and orient it according to a random topological ordering. We then generate a feedforward neural network $f_{k}$ with a single hidden layer of 10 neurons for $k=1,\dots,p$. The weights of $f_{k}$ are sampled uniformly from $[-0.7,-0.3]\cup[0.3,0.7]$, with hidden layer weights of node $k$’s non-parents set to zero. The activation is a ReLU. The noise $\varepsilon_{1},\dots,\varepsilon_{n}$ is drawn iid and the variables are taken as
@@
Appendix G Additional experiments
Figure 6: Performance on synthetic datasets generated from linear Erdős–Rényi DAGs with $p=20$ nodes, $s=60$ edges, and Gaussian noise. The averages (solid points) and standard errors (error bars) are measured over 10 independently and identically generated datasets.
@@ Appendix G Additional experiments
Figure 9: Performance on synthetic datasets generated from linear Erdős–Rényi DAGs with $p=20$ nodes, $s=40$ edges, and exponential noise. The averages (solid points) and standard errors (error bars) are measured over 10 independently and identically generated datasets.
Figure 10: Performance on synthetic datasets generated from linear scale-free (degree exponent 2) DAGs with $p=20$ nodes, $s=40$ edges, and Gaussian noise. The averages (solid points) and standard errors (error bars) are measured over 10 independently and identically generated datasets.
Figure 11: Performance on synthetic datasets generated from linear scale-free (degree exponent 3) DAGs with $p=20$ nodes, $s=40$ edges, and Gaussian noise. The averages (solid points) and standard errors (error bars) are measured over 10 independently and identically generated datasets.
We investigate the robustness of ProDAG to several alternative simulation designs. Figures 6 and 7 report results on datasets simulated from dense graphs containing two and three times as many edges as used in the main experiments. Figure 8 and 9 report the results from non-Gaussian noise distributions, specifically Gumbel and exponential distributions. Finally, Figure 10 and 11 report results on scale-free graphs instead of Erdős–Rényi graphs, where the degree exponent is set to 2 and 3. Under all three simulation designs, the relative performance of ProDAG is roughly similar to that in the main experiments, indicating that ProDAG is reliable across a broad range of realistic scenarios.
Appendix H Limitations
Though our approach has several favorable properties, it naturally has limitations, the main ones we highlight here. As with most Bayesian approaches that employ variational inference, ours does not guarantee that the learned variational posterior approximates the true posterior well. The possibility of a poor variational approximation is further complicated by the nonconvexity of the acyclicity constraint, giving rise to a nonconvex optimization problem that can be solved only to a stationary point. We remark, however, that nonconvexity is a property of most existing DAG learning methods.
Appendix I Societal impacts
DAG learning tools play a key role in causal inference, where the goal is to establish causal relationships between random variables. Our approach benefits from quantifying uncertainty in this process, which can help guide downstream decision-making. However, as with any DAG learning tool applied to observational data, one cannot causally interpret the DAGs learned by our approach without making empirically unverifiable assumptions such as causal sufficiency. Therefore, it is crucial to involve subject-matter experts capable of determining the validity of any such assumptions.